# Global Buckley–Leverett for Multicomponent Flow in Fractured Media: Isothermal Equation-of-State Coupling and Dynamic Capillarity

### v1

@@ II.3 Momentum with non-Darcy and dynamic capillarity
$-\nabla\Big(p-\rho_{\alpha}^{\kappa}\mathbf{g}\!\cdot\!\mathbf{x}\Big)\;-\;\nabla p_{c,\alpha}^{\kappa,\mathrm{dyn}}(S_{\alpha}^{\kappa})$
$=\frac{\mu_{\alpha}^{\kappa}}{k^{\kappa}(p,\sigma^{\prime})\,k_{r\alpha}^{\kappa}(S)}\,\mathbf{v}_{\alpha}^{\kappa}\;+\;\beta_{\alpha}^{\kappa}\,\rho_{\alpha}^{\kappa}\,\|\mathbf{v}_{\alpha}^{\kappa}\|\,\mathbf{v}_{\alpha}^{\kappa},$ (23)
where the dynamic capillary pressure is split into a quasi-static (equilibrium) part and a rate correction,
$p_{c,\alpha}^{\kappa,\mathrm{dyn}}(S_{\alpha}^{\kappa})$ $=p_{c,\alpha}^{\mathrm{AM}}(S_{\alpha}^{\kappa};\,\gamma_{\alpha},\theta_{\alpha},\mathcal{M}_{\kappa},\xi_{\kappa})$
$-\tau_{\alpha}^{\kappa}(S_{\alpha}^{\kappa})\,\partial_{t}S_{\alpha}^{\kappa}.$ (24)
The Alonso–Marroquín pore-morphology [AlonsoMarroquinAndersson2025PMM] method maps saturation to an effective capillary radius selected by invasion under Young–Laplace balance. A quasi-2D curvature gives
@@
Two routes are practical: (i) image-informed—infer the throat-radius CDF $F_{r}(r)$ from micro-CT or mercury intrusion, define an effective saturation $S_{e}$ (accounting for residuals and trapping $\xi_{\kappa}$), set $r_{c}(S)=F_{r}^{-1}(1-S_{e})$, and tabulate $p_{c,\alpha}^{\mathrm{AM}}(S)$ from (26); (ii) surrogate fit—when detailed morphology is unavailable, fit a monotone surrogate (e.g., Brooks–Corey, van Genuchten, or a smooth spline) to measurements and interpret it as $p_{c,\alpha}^{\mathrm{AM}}(S)$. The continuum model needs only a smooth $p_{c}(S)$; the morphology furnishes a constructive path and hysteresis via $\xi_{\kappa}$.
@@
In fractures, inertial departures from Darcy are not optional: rough-walled data show measurable non-Darcy losses at modest in-plane Reynolds numbers, with losses increasing as aperture decreases [Chen2015JHydrol]. We therefore take $\beta_{\alpha}^{\mathrm{f}}>0$ by default (calibrated from fracture tests) and $\beta_{\alpha}^{\mathrm{m}}=0$ in matrix unless core data indicate otherwise. A convenient dimensionless trigger for significance is
@@ II.4 Buckley–Leverett–style fractional–flow split
To avoid a clash with the fugacity notation $f_{i}^{\alpha}$ used earlier, we denote the fractional–flow weights by
@@
Thus Forchheimer effects enter as colinearity damping folded into $\tilde{\lambda}_{\alpha}^{\kappa}$. No global pressure is assumed in this derivation; when the TD/gTD condition holds, the same split coexists with a cleaner pressure equation (see §II.5).
@@ II.5 Global pressure and the generalized TD hypothesis
is exact on the ternary simplex, i.e. the TD compatibility [Chavent2009, DiChiara2010JCP]:
@@ II.6 Classical limit and well-posedness
Work in a fixed continuum $\kappa$ and suppress κ for readability. Impose the standard BL hypotheses: (a) incompressible rock and fluids ($\phi$ and $\rho_{\alpha}$ constants); (b) fixed phase compositions and no interphase mass exchange (each $x_{i\alpha}$ constant, hence each phase molar density $c_{\alpha}$ constant up to $S_{\alpha}$ factors); (c) no multicomponent diffusion/dispersion ($\mathbf{J}_{i\alpha}\equiv\mathbf{0}$ in (6)–(9)); (d) negligible capillarity ($p_{c,\alpha}^{\mathrm{AM}}\equiv 0$ and $\tau_{\alpha}\equiv 0$ in (24)); (e) Darcy momentum (set $\beta_{\alpha}=0$ in (23)); (f) constant permeability $k\equiv k_{0}$ (freeze (11)). Aggregating the conservative balances (14) over the components that reside in phase $\alpha$ (and defining a phase source $q_{\alpha}$ consistent with the fixed-composition limit) gives the phase saturation laws
@@
$-\sum_{\alpha}\int_{\Omega}c_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}^{\!\top}\,\mathbf{D}_{\alpha}^{\mathrm{MS}}\,\boldsymbol{\Gamma}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}\,dx\;\leq\;0,$ (62)
with $\boldsymbol{\Gamma}_{\alpha}=\partial\boldsymbol{\mu}_{\alpha}/\partial(\ln\boldsymbol{x}_{\alpha})$ SPD in stable phases. Control of $\nabla\boldsymbol{x}_{\alpha}$ lifts, via the EOS map (2)–(4), to control of $\nabla z$.
@@
$\frac{d}{dt}\,\Big[\mathcal{F}(z)$ $+\mathcal{E}_{c}(S)\Big]+\underbrace{\sum_{\alpha}\!\int_{\Omega}c_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}^{\!\top}\,\mathbf{D}_{\alpha}^{\mathrm{MS}}\,\boldsymbol{\Gamma}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}\,dx}_{\text{controls }\nabla z}$
@@ II.7 Fracture modeling and non-Darcy effects
(i) Dual–continuum: matrix–fracture coupling appears as conservative sources $T_{i}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ assembled via (17)–(19). (ii) EDFM/pEDFM: fractures are lower-dimensional control volumes coupled to matrix cells by NNCs using (20)–(22) [WarrenRoot1963, Olorode2020pEDFM].
Let $b$ be the hydraulic aperture and $\zeta_{r}\!\in\!(0,1]$ a roughness factor. The in-plane permeability and transmissivity are
@@
where $\chi_{\alpha}^{\mathrm{f}}$ is given by (39) after the substitutions $\{k,k_{r\alpha},\mu_{\alpha},\rho_{\alpha},u_{\alpha}^{D}\}\to\{k^{\mathrm{f}},k_{r\alpha}^{\mathrm{f}},\mu_{\alpha}^{\mathrm{f}},\rho_{\alpha}^{\mathrm{f}},u_{\alpha}^{\mathrm{f},D}\}$. For weak inertia, use the linearization (40). A convenient activation metric is the in-plane Reynolds number $\mathrm{Re}_{\tau}^{\alpha}=(\rho_{\alpha}^{\mathrm{f}}\|\mathbf{v}_{\alpha}^{\mathrm{f},D}\|\,b)/\mu_{\alpha}^{\mathrm{f}}$; rough-walled data show non-Darcy behavior at modest $\mathrm{Re}_{\tau}^{\alpha}$ and increasing losses as $b$ decreases [Chen2015JHydrol].

### v2

@@ II.3 Momentum with non-Darcy and dynamic capillarity
$-\nabla\!\Big(p-\rho_{\alpha}^{\kappa}\,\mathbf{g}\!\cdot\!\mathbf{x}\Big)$
$-\nabla p_{c,\alpha}^{\kappa,\mathrm{dyn}}(S_{\alpha}^{\kappa})=\frac{\mu_{\alpha}^{\kappa}}{k^{\kappa}(p,\sigma^{\prime})\,k_{r\alpha}^{\kappa}(S)}\,\mathbf{v}_{\alpha}^{\kappa}+\beta_{\alpha}^{\kappa}\,\rho_{\alpha}^{\kappa}\,\|\mathbf{v}_{\alpha}^{\kappa}\|\,\mathbf{v}_{\alpha}^{\kappa},$ (22)
where $k^{\kappa}(p,\sigma^{\prime})$ is the stress-sensitive intrinsic permeability and $k_{r\alpha}^{\kappa}(S)$ is the relative permeability. The Forchheimer coefficient $\beta_{\alpha}^{\kappa}$ is treated as a calibrated closure: in fractures it may depend on geometry (hydraulic aperture and roughness) and thus can be stress dependent through $b(\sigma^{\prime})$, whereas in the matrix we typically set $\beta_{\alpha}^{\mathrm{m}}=0$ unless core data indicate otherwise. The Forchheimer term here is used as a pragmatic inertial damping for high-rate fracture flow; it does not represent interface-mobilization-induced nonlinear multiphase rheology (often manifesting as power-law Darcy-scale behavior) [15]. When needed, the framework admits alternative nonlinear Darcy-scale closures by replacing the per-phase momentum relation with a generalized mobility operator, e.g., $\mathbf{v}_{\alpha}^{\kappa}=-\,k^{\kappa}\,\lambda_{\alpha}^{\kappa}\,\mathcal{M}_{\alpha}\!\big(|\nabla\Phi_{\alpha}^{\kappa}|\big)\,\nabla\Phi_{\alpha}^{\kappa},$ without changing the conservative component balances or the global-pressure/fractional-flow structure [15, 38].
The dynamic capillary pressure is decomposed into an equilibrium contribution and a rate-dependent correction,
$p_{c,\alpha}^{\kappa,\mathrm{dyn}}(S_{\alpha}^{\kappa})$ $=p_{c,\alpha}^{\mathrm{AM}}(S_{\alpha}^{\kappa};\gamma_{\alpha},\theta_{\alpha},\mathcal{M}_{\kappa},\xi_{\kappa})-\tau_{\alpha}^{\kappa}(S_{\alpha}^{\kappa})\,\partial_{t}S_{\alpha}^{\kappa}.$ (23)
The Alonso–Marroquín & Andersson pore-morphology [2] method maps saturation to an effective capillary radius selected by invasion under Young–Laplace balance. A quasi-2D curvature gives
@@
Two routes are practical: (i) image-informed—infer the throat-radius cumulative distribution function (CDF) $F_{r}(r)$ from micro-CT or mercury intrusion, define an effective saturation $S_{e}$ (accounting for residuals and trapping $\xi_{\kappa}$), set $r_{c}(S)=F_{r}^{-1}(1-S_{e})$, and tabulate $p_{c,\alpha}^{\mathrm{AM}}(S)$ from (25); (ii) surrogate fit—when detailed morphology is unavailable, fit a monotone surrogate (e.g., Brooks–Corey, van Genuchten, or a smooth spline) to measurements and interpret it as $p_{c,\alpha}^{\mathrm{AM}}(S)$. The continuum model needs only a smooth $p_{c}(S)$; the morphology furnishes a constructive path and hysteresis via $\xi_{\kappa}$.
@@
In fractures, inertial departures from Darcy are not optional: rough-walled data show measurable non-Darcy losses at modest in-plane Reynolds numbers, with losses increasing as aperture decreases [11]. We therefore take $\beta_{\alpha}^{\mathrm{f}}>0$ by default (calibrated from fracture tests) and $\beta_{\alpha}^{\mathrm{m}}=0$ in matrix unless core data indicate otherwise. A convenient dimensionless trigger for significance is
@@ II.4 Buckley–Leverett–style fractional–flow split
Here, $\mathbf{v}_{\alpha}$ denotes the phase superficial (Darcy) velocity (volumetric flux per bulk cross-sectional area) [6, 34]. The total superficial velocity is defined as $\mathbf{v}_{t}=\sum_{\alpha}\mathbf{v}_{\alpha}$ [29, 5]. Pore (seepage) velocities are obtained by dividing by $\phi S_{\alpha}$ if needed [6]. To avoid a clash with the fugacity notation $f_{i}^{\alpha}$ used earlier, we denote the fractional–flow weights by
@@
An overbar $\bar{(\cdot)}_{\lambda}$ denotes a mobility-weighted average using $\lambda_{\alpha}$; a tilde $\tilde{(\cdot)}$ denotes Forchheimer-damped (‘apparent’) quantities using $\tilde{\lambda}_{\alpha}=\chi_{\alpha}\lambda_{\alpha}$; superscript D denotes the Darcy-level (no inertia) flux. Thus Forchheimer effects enter as colinearity damping folded into $\tilde{\lambda}_{\alpha}^{\kappa}$. No global pressure is assumed in this derivation; when the TD/gTD condition holds, the same split coexists with a cleaner pressure equation (see §II.5).
@@ II.5 Global pressure and the generalized TD hypothesis
is exact on the ternary simplex, i.e. the TD compatibility [9, 13]:
@@ II.6 Classical limit and well-posedness
A fixed continuum $\kappa$ is assumed, and κ is suppressed for readability. The standard BL hypotheses are imposed: (a) incompressible rock and fluids ($\phi$ and $\rho_{\alpha}$ constants); (b) fixed phase compositions and no interphase mass exchange (each $x_{i\alpha}$ constant, hence each phase molar density $c_{\alpha}$ constant up to $S_{\alpha}$ factors); (c) no multicomponent diffusion/dispersion ($\mathbf{J}_{i\alpha}\equiv\mathbf{0}$ in (6)–(8)); (d) negligible capillarity ($p_{c,\alpha}^{\mathrm{AM}}\equiv 0$ and $\tau_{\alpha}\equiv 0$ in (23)); (e) Darcy momentum (set $\beta_{\alpha}=0$ in (22)); (f) constant permeability $k\equiv k_{0}$ (freeze (11)). Aggregating the conservative balances (14) over the components that reside in phase $\alpha$ (and defining a phase source $q_{\alpha}$ consistent with the fixed-composition limit) gives the phase saturation laws
@@
$-\sum_{\alpha}\int_{\Omega}c_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}^{\!\top}\,\mathbf{D}_{\alpha}^{\mathrm{MS}}\,\boldsymbol{\mathcal{G}}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}\,dx\;\leq\;0,$ (61)
with $\boldsymbol{\mathcal{G}}_{\alpha}=\partial\boldsymbol{\mu}_{\alpha}/\partial(\ln\boldsymbol{x}_{\alpha})$ SPD in stable phases. Control of $\nabla\boldsymbol{x}_{\alpha}$ lifts, via the EOS map (2)–(4), to control of $\nabla z$.
@@
$\frac{d}{dt}\,\Big[\mathcal{F}(z)$ $+\mathcal{E}_{c}(S)\Big]+\underbrace{\sum_{\alpha}\!\int_{\Omega}c_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}^{\!\top}\,\mathbf{D}_{\alpha}^{\mathrm{MS}}\,\boldsymbol{\mathcal{G}}_{\alpha}\,\nabla\boldsymbol{x}_{\alpha}\,dx}_{\text{controls }\nabla z}$
@@ II.7 Fracture modeling and non-Darcy effects
(i) Dual–continuum: matrix–fracture coupling appears as conservative sources $T_{i}^{\mathrm{m}\leftrightarrow\mathrm{f}}$ assembled via (17)–(18). (ii) EDFM/pEDFM: fractures are lower-dimensional control volumes coupled to matrix cells by NNCs using (19)–(21) [36, 28].
Let $b$ be the hydraulic aperture and $\zeta_{r}\!\in\!(0,1]$ a conductivity reduction factor: $\zeta_{r}=1$ gives the smooth parallel-plate cubic law, while smaller values represent roughness/contact-induced reductions; $\zeta_{r}\to 0$ corresponds to an effectively closed/nonconductive pathway. The in-plane permeability and transmissivity are
@@
where $\chi_{\alpha}^{\mathrm{f}}$ is given by (38) after the substitutions $\{k,k_{r\alpha},\mu_{\alpha},\rho_{\alpha},u_{\alpha}^{D}\}\to\{k^{\mathrm{f}},k_{r\alpha}^{\mathrm{f}},\mu_{\alpha}^{\mathrm{f}},\rho_{\alpha}^{\mathrm{f}},u_{\alpha}^{\mathrm{f},D}\}$. For weak inertia, use the linearization (39). A convenient activation metric is the in-plane Reynolds number $\mathrm{Re}_{\tau}^{\alpha}=(\rho_{\alpha}^{\mathrm{f}}\|\mathbf{v}_{\alpha}^{\mathrm{f},D}\|\,b)/\mu_{\alpha}^{\mathrm{f}}.$ Rough-walled fracture studies report measurable departures from Darcy/cubic-law behavior once $\mathrm{Re}_{\tau}^{\alpha}=\mathcal{O}(1)$, while Forchheimer-type quadratic corrections often provide good fits in the stronger-inertia regime $\mathrm{Re}_{\tau}^{\alpha}\gtrsim 20$ (geometry/roughness dependent) [1, 37]. Consistent with these observations, increasing roughness and decreasing hydraulic aperture tend to enhance inertial losses; accordingly, the Forchheimer coefficient is not universal in fractures and is commonly modeled as $\beta_{\alpha}^{\mathrm{f}}=\beta_{\alpha}^{\mathrm{f}}(b,\mathrm{roughness})$ with $b=b(\sigma^{\prime})$, with $\beta_{\alpha}^{\mathrm{f}}$ typically increasing as aperture decreases (often approximated by inverse-aperture trends over limited ranges) [11].
@@ Conclusions
Notation and acronyms
Acronyms
BL
Buckley–Leverett.
EOS
Equation of state.
PVT
Pressure–volume–temperature properties.
GBL-NN
Global Buckley–Leverett for $N_{c}$ components and $N_{p}$ phases.
MS
Maxwell–Stefan (multicomponent diffusion).
TD/gTD
Total-differential / generalized total-differential compatibility condition for global pressure.
EDFM
Embedded discrete fracture model.
pEDFM
Projection-based embedded discrete fracture model.
NNC
Non-neighbor connection (matrix–fracture coupling).
SRK
Soave–Redlich–Kwong cubic EOS.
PR
Peng–Robinson cubic EOS.
CDF
Cumulative distribution function.
Symbols (primary fields, thermodynamics, and transport)
$\kappa\in\{\mathrm{m},\mathrm{f}\}$
Continuum index: matrix ($\mathrm{m}$) and fracture ($\mathrm{f}$).
$\alpha=1,\dots,N_{p}$
Phase index; $N_{p}^{\star}\leq N_{p}$ phases may be present at equilibrium.
$i=1,\dots,N_{c}$
Component index.
$\mathbf{x}$
Spatial position vector.
$t$
Time.
$T$
Temperature (fixed; isothermal).
$p^{\kappa}$
Thermodynamic pressure used in EOS flash (per continuum).
$p_{g}^{\kappa}$
Global pressure (mobility-weighted scalar driving Darcy total flux in a continuum).
$S_{\alpha}^{\kappa}$
Phase saturation in continuum $\kappa$, with $\sum_{\alpha}S_{\alpha}^{\kappa}=1$.
$z_{i}$
Overall (bulk) mole fraction, with $\sum_{i}z_{i}=1$.
$\nu_{\alpha}$
Phase molar fraction from EOS flash, with $\sum_{\alpha}\nu_{\alpha}=1$.
$x_{i\alpha}$
Mole fraction of component $i$ in phase $\alpha$, with $\sum_{i}x_{i\alpha}=1$.
$M_{i}$
Molar mass of component $i$.
$\bar{M}_{\alpha}$
Phase mixture molar mass, $\bar{M}_{\alpha}=\sum_{i}M_{i}x_{i\alpha}$.
$c_{\alpha}$
Phase molar density.
$\rho_{\alpha}$
Phase mass density, $\rho_{\alpha}=\bar{M}_{\alpha}c_{\alpha}$.
$\mu_{\alpha}$
Phase dynamic viscosity (from transport-property correlations evaluated at flashed state).
$\mu_{i\alpha}$
Chemical potential of component $i$ in phase $\alpha$.
$f_{i}^{\alpha}$
Fugacity of component $i$ in phase $\alpha$; $\varphi_{i\alpha}$ is the fugacity coefficient.
Fluxes, mobilities, and averaging conventions
$\mathbf{v}_{\alpha}^{\kappa}$
Phase superficial (Darcy) velocity in continuum $\kappa$ (volumetric flux per bulk cross-sectional area).
$\mathbf{v}_{t}^{\kappa}$
Total superficial velocity, $\mathbf{v}_{t}^{\kappa}=\sum_{\alpha}\mathbf{v}_{\alpha}^{\kappa}$.
$\mathbf{v}_{\alpha}^{\kappa,D}$
Darcy-level (no inertia) phase velocity.
$\mathbf{v}_{t}^{\kappa,D}$
Darcy-level total velocity.
$k^{\kappa}$
Intrinsic permeability (stress sensitive).
$k_{r\alpha}^{\kappa}(S)$
Relative permeability of phase $\alpha$.
$\lambda_{\alpha}^{\kappa}$
Phase mobility, $\lambda_{\alpha}^{\kappa}=k_{r\alpha}^{\kappa}/\mu_{\alpha}^{\kappa}$.
$\Lambda_{t}^{\kappa}$
Total mobility, $\Lambda_{t}^{\kappa}=\sum_{\alpha}\lambda_{\alpha}^{\kappa}$.
$F_{\alpha}^{\kappa}$
Fractional-flow weight, $F_{\alpha}^{\kappa}=\lambda_{\alpha}^{\kappa}/\Lambda_{t}^{\kappa}$.
$\chi_{\alpha}^{\kappa}$
Forchheimer damping factor (per phase), $0<\chi_{\alpha}^{\kappa}\leq 1$.
$\tilde{\lambda}_{\alpha}^{\kappa}$
Apparent (damped) mobility, $\tilde{\lambda}_{\alpha}^{\kappa}=\chi_{\alpha}^{\kappa}\lambda_{\alpha}^{\kappa}$.
$\tilde{F}_{\alpha}^{\kappa}$
Apparent fractional-flow weight, $\tilde{F}_{\alpha}^{\kappa}=\tilde{\lambda}_{\alpha}^{\kappa}/\tilde{\Lambda}_{t}^{\kappa}$.
$\bar{(\cdot)}_{\lambda}$
Mobility-weighted average using $\lambda_{\alpha}$ (e.g., $\bar{\rho}_{\lambda}=\sum_{\alpha}\lambda_{\alpha}\rho_{\alpha}/\Lambda_{t}$).
$\tilde{(\cdot)}$
Quantity formed with apparent mobilities $\tilde{\lambda}_{\alpha}$ (e.g., $\tilde{\bar{\rho}}_{\lambda}$).
Diffusion, capillarity, and adsorption
$\mathbf{J}_{i\alpha}$
Maxwell–Stefan diffusive molar flux of component $i$ in phase $\alpha$ (relative to phase-average velocity), satisfying $\sum_{i}\mathbf{J}_{i\alpha}=\mathbf{0}$.
$\mathbf{D}^{\mathrm{MS}}_{ij,\alpha}$
Effective Maxwell–Stefan diffusivity (isotropic scalar times $\mathbf{I}$ or anisotropic spatial tensor).
$\mathbf{D}^{\mathrm{MS}}_{\alpha}$
Compact operator form of MS diffusivity for phase $\alpha$.
$\boldsymbol{\mathcal{G}}_{\alpha}$
Thermodynamic-factor matrix in MS diffusion, $\boldsymbol{\mathcal{G}}_{\alpha}]_{ij}=\partial\mu_{i\alpha}/\partial\ln x_{j\alpha}$.
$p_{c,\alpha}^{\kappa,\mathrm{dyn}}$
Dynamic capillary pressure (per phase, per continuum).
$p_{c,\alpha}^{\mathrm{AM}}(S)$
Equilibrium capillary pressure from the pore-morphology (Alonso–Marroquín) model or a monotone surrogate fit.
$\tau_{\alpha}^{\kappa}(S)$
[…]
$\Gamma_{i}^{\kappa}$
Adsorbed inventory (storage term) of component $i$ in continuum $\kappa$.
$t_{i}^{\kappa}$
Adsorption relaxation time (units s).
Fractures and matrix–fracture transfer
$b$
Hydraulic aperture of a fracture.
$\zeta_{r}$
Conductivity reduction factor relative to smooth parallel-plate cubic law, $\zeta_{r}\in(0,1]$.
$k^{\mathrm{f}}$
Fracture in-plane permeability, $k^{\mathrm{f}}=\zeta_{r}b^{2}/12$.
$\mathcal{T}^{\mathrm{f}}$
Fracture transmissivity, $\mathcal{T}^{\mathrm{f}}=\zeta_{r}b^{3}/12$.
$\beta_{\alpha}^{\kappa}$
Forchheimer coefficient (calibrated; fracture values may depend on $b$ and roughness).
$\mathrm{Re}_{\tau}^{\alpha}$
Fracture-plane Reynolds number, $\mathrm{Re}_{\tau}^{\alpha}=(\rho_{\alpha}^{\mathrm{f}}\|\mathbf{v}_{\alpha}^{\mathrm{f},D}\|\,b)/\mu_{\alpha}^{\mathrm{f}}$.
$\Phi_{\alpha}^{\kappa}$
Phase hydraulic potential, $\Phi_{\alpha}^{\kappa}=p^{\kappa}-p_{c,\alpha}^{\kappa,\mathrm{dyn}}-\rho_{\alpha}^{\kappa}\mathbf{g}\!\cdot\!\mathbf{x}$.
$\omega$
Dual-porosity shape factor (units 1/m2).
$k_{\mathrm{int}}$
Interporosity (matrix–fracture) permeability scale.
$T_{mf}$
Geometric transmissibility for matrix–fracture NNC.
$\nabla_{\tau}$
Tangential (fracture-plane) gradient operator.